OpenStudy (anonymous):

) Find the slope m of the tangent to the curve y = 7/radical x at the point where x = a > 0.

OpenStudy (anonymous):

y= \[7\div \sqrt{x}\]

jimthompson5910 (jim_thompson5910):

\[\Large y = \frac{7}{\sqrt{x}}\] \[\Large y^{\prime} = \frac{d}{dx}\left(\frac{7}{\sqrt{x}}\right)\] \[\Large y^{\prime} = \frac{d}{dx}\left(7x^{-\frac{1}{2}}\right)\] \[\Large y^{\prime} = 7\frac{d}{dx}\left(x^{-\frac{1}{2}}\right)\] \[\Large y^{\prime} = 7(-\frac{1}{2})x^{-\frac{3}{2}}\] \[\Large y^{\prime} = -\frac{7}{2x^{\frac{3}{2}}}\] \[\Large y^{\prime} = -\frac{7}{2x\sqrt{x}}\]

OpenStudy (anonymous):

That answer was not taken by the website?

jimthompson5910 (jim_thompson5910):

that's just the derivative function

jimthompson5910 (jim_thompson5910):

you use that to find the slope at the given point on the function

OpenStudy (anonymous):

oh, it says at the point where x=a>0

jimthompson5910 (jim_thompson5910):

in your case, x = a, so replace each x with a to get \[\Large y^{\prime} = -\frac{7}{2x\sqrt{x}}\] \[\Large y^{\prime} = -\frac{7}{2a\sqrt{a}}\] So \[\Large m = -\frac{7}{2a\sqrt{a}}\]

jimthompson5910 (jim_thompson5910):

at x = a

OpenStudy (anonymous):

ok, how do i find equations of the tangent lines at the points (1, 7) and (4, 7/2)?

OpenStudy (anonymous):

two separate equations for each point..

jimthompson5910 (jim_thompson5910):

same function?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

plug in a = 1 (for the first point) to get the slope \[\Large m = -\frac{7}{2a\sqrt{a}}\] \[\Large m = -\frac{7}{2(1)\sqrt{1}}\] \[\Large m = -\frac{7}{2}\] Now use the general line equation y = mx+b along with x = 1 and x = 7 to get y = mx+b 7 = (-7/2)(1)+b 7 = -7/2 + b 7+7/2 = b 21/2 = b So the equation of the tangent line at the point (1,7) is \[\Large y = -\frac{7}{2}x+\frac{21}{2}\] So the same with the second point

OpenStudy (anonymous):

okay, thank you!